# Discovering environments with XRM

### v1

@@ 6 Discussion
We have introduced Cross-Risk Minimization (XRM), a simple algorithm for environment discovery. XRM provides a recipe to tune its hyper-parameters, does not require early-stopping, and can discover environments for all training and validation data—dropping the requirement for human annotations at all. More specifically, XRM trains two twin label predictors on random halves of the training data, while encouraging each twin to imitate confident held-out mistakes by their sibling. This implements an “echo-chamber” that identifies environments that differ in spurious correlation, and endow domain generalization algorithms with oracle-like performance.
We highlight two directions for future work. Firstly, how does XRM relate to the invariance principle $Y\perp E\mid\Phi(X)$? What is the interplay between revealing relevant labels $Y$ and relevant environments $E$ as to afford invariance? To our knowledge, XRM is the first environment discovery algorithm tampering with labels $Y$, thus exploring invariance—and the violation thereof—from a new angle. Because relabeling happens with a probability proportional to confidence, we expect model calibration to play a role in understanding the theoretical underpinnings of XRM, as it happened with other invariance methods (Wald et al., 2021). Overall, the theoretical analysis of XRM will call for new tools, because the optimal solution is a moving target that depends on the dynamic evolution of the labels, steering away from the Bayes-optimal predictor $P(Y\mid X)$.
Secondly, we would like to further understand the relationship between XRM and the multifarious phenomenon of memorization. Good memorization affords invariance (Where did I park my car?), and therefore depends on the collection of environments deemed relevant. Bad memorization happens due to “structured over-fitting”, commonly incarnated as a bad learning strategy “use a simple feature for the majority, then memorize the minority”. XRM seems to attack a similar problem but, how does it specifically relate to these two flavours of memorization? Does XRM discover environments that promote features that benefit all examples?
@@ Acknowledgements
We thank Kartik Ahuja, Martín Arjovsky, Léon Bottou, Kamalika Chaudhuri, Cijo Jose, Reyhane Askari Hemmat, Polina Kirichenko and others for fruitful discussions. We thank the FAIR leadership for their help in publishing this manuscript.
Appendix
Appendix A Experimental details
We follow the experimental protocol of SubpopBench [Yang et al., 2023]. Therefore, image datasets use a pretrained ResNet-50 [He et al., 2016] unless otherwise mentioned, and text datasets use a pretrained BERT [Devlin et al., 2018]. All images are resized and center-cropped to $224\times 224$ pixels, and undergo no data augmentation. We use SGD with momentum $0.9$ to learn from image datasets unless otherwise mentioned, and we employ AdamW [Loshchilov and Hutter, 2017] with default $\beta_{1}=0.9$ and $\beta_{2}=0.999$ for text benchmarks. For the ColorMNIST experiment [Arjovsky et al., 2019], we train a three-layer fully-connected network with layer sizes $[2*14*14,300,300,2]$ and use ReLU as the activation function. The network is optimized using the Adam optimizer with a learning rate of $31e-$ 3, and default parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. For the experiment on CIFAR-10 [Krizhevsky et al., 2009], we train a VGG-16 model [Simonyan and Zisserman, 2014] using SGD with a learning rate of $21e-$ 2 and a momentum of $0.9$ We train XRM and phase-2 algorithms for a number of iterations that allows convergence within a reasonable compute budget. These are 5,000 steps for Waterbirds and Metashift, 10,000 steps for ImageNetBG, 20,000 steps for MultiNLI, and 30,000 steps for CivilComments.
A.1 Phase-1 XRM model selection
For phase-1, we run XRM with 16 random combinations of hyper-parameters, each over $k_{1}=10$ random seeds. (We repeat runs with null accuracy on one of the classes.) For each of the 16 hyper-parameter combinations, we average the number of flipped labels appearing at the last iteration (early-stopping is not necessary with XRM) across the 10 seeds. This will tell us which one of the 16 hyper-parameter combinations is best. For that combination, we average the training and validation logit matrices across the 10 random seeds. Finally, we discover environments using equation 4.
A.2 Phase-2 DG model selection.
For all phase-2 domain generalization algorithms (ERM, SUBG, RWG, GroupDRO), we search over 16 random combinations of hyper-parameters. We select the hyper-parameter combination and early-stopping iteration yielding maximal worst-group-accuracy (or, in the absence of groups, worst-class-accuracy). We re-run the best hyper-parameter combination for $k_{2}=10$ random seeds to report avg/std for test worst-group-accuracies (always computed with respect to the ground-truth group annotations). The $k_{1}$ random seeds from phase-1 do not contribute to error bars.
A.3 Hyper-parameter sampling grids
algorithm hyper-parameter ResNet BERT
learning rate $10^{\text{Uniform}(-4,-2)}$ $10^{\text{Uniform}(-5.5,-4)}$
SUBG, RWG batch size $2^{\text{Uniform}(6,7)}$ $2^{\text{Uniform}(3,5.5)}$
dropout $\text{Random}([0,0.1,0.5])$
GroupDRO $\eta$ $10^{\text{Uniform}(-3,-1)}$ $10^{\text{Uniform}(-3,-1)}$
Appendix B Learning to Split on Waterbirds
We benchmarked the official learning to split code-base https://github.com/YujiaBao/ls on the WaterBirds dataset. We assessed the method’s sensitivity to two hyperparameters: the number of epochs used for early stopping (patience argument in the codebase) and the pre-supposed ratio of groups (based on the ratio argument in the code). For patience we swept over (2, 5, 10) with 5 being the default value. For ratio, we swept over (0.25, 0.5, 0.75) with 0.75 being the default value based on the paper. We found worst group performance using a fixed GroupDRO phase-2 training varied by as much as ±$\pm 7\%{}$ on Waterbirds.
Appendix C Experimental results with error bars
@@
Waterbirds 70.4 $\pm$2.99 76.1 $\pm$2.37 75.3 $\pm$1.96 71.7 $\pm$4.09 88.0 $\pm$2.61 86.1 $\pm$1.28 74.8 $\pm$2.50 87.0 $\pm$1.63 84.5 $\pm$1.53 73.0 $\pm$2.75 86.7 $\pm$1.00 76.3 $\pm$8.41
CelebA 62.7 $\pm$2.73 71.9 $\pm$3.48 67.6 $\pm$3.48 68.8 $\pm$1.29 89.1 $\pm$1.67 89.8 $\pm$1.39 70.7 $\pm$1.32 89.5 $\pm$1.45 88.0 $\pm$2.56 68.5 $\pm$2.13 87.1 $\pm$2.70 87.5 $\pm$2.54
MultiNLI 54.8 $\pm$4.04 65.3 $\pm$3.02 65.8 $\pm$3.41 69.7 $\pm$2.65 76.1 $\pm$1.29 74.3 $\pm$1.88 69.3 $\pm$1.77 71.1 $\pm$1.60 73.3 $\pm$1.56 53.7 $\pm$2.97 72.8 $\pm$0.66 71.3 $\pm$1.58
CivilComments 55.1 $\pm$3.46 59.7 $\pm$5.77 61.4 $\pm$4.48 59.3 $\pm$2.05 69.3 $\pm$2.32 71.3 $\pm$1.35 54.0 $\pm$4.58 65.8 $\pm$6.30 73.4 $\pm$0.93 58.7 $\pm$2.56 64.0 $\pm$7.63 72.9 $\pm$1.12
ColorMNIST 10.1 $\pm$0.51 10.1 $\pm$2.40 26.9 $\pm$2.27 10.0 $\pm$0.51 10.1 $\pm$2.37 70.5 $\pm$0.98 10.1 $\pm$0.51 10.2 $\pm$1.85 70.2 $\pm$1.00 10.1 $\pm$0.51 10.0 $\pm$2.21 70.8 $\pm$1.09
MetaShift 70.8 $\pm$2.99 67.7 $\pm$4.62 71.2 $\pm$3.95 70.8 $\pm$3.91 73.7 $\pm$4.42 71.2 $\pm$4.81 66.5 $\pm$4.55 70.8 $\pm$4.45 69.7 $\pm$4.86 70.0 $\pm$3.38 73.5 $\pm$3.49 69.2 $\pm$5.58
ImagenetBG 75.6 $\pm$3.04 76.9 $\pm$1.89 76.9 $\pm$1.93 73.0 $\pm$3.43 76.1 $\pm$1.40 75.9 $\pm$1.69 76.9 $\pm$2.42 76.5 $\pm$2.65 77.0 $\pm$2.66 75.0 $\pm$3.55 75.0 $\pm$3.55 76.4 $\pm$1.79
Appendix D XRM in PyTorch

### v2

@@ 7 Discussion
We have introduced Cross-Risk Minimization (XRM), a simple algorithm for environment discovery. XRM provides a recipe to tune its hyper-parameters, does not require early-stopping, and can discover environments for all training and validation data—dropping the requirement for human annotations at all. More specifically, XRM trains two twin classifiers on random halves of the training data, while encouraging each twin to imitate confident held-out mistakes by their sibling. This implements an “echo-chamber” that identifies environments that differ only in spurious correlation, and endow subsequent OOD generalization algorithms with oracle-like performance.
We highlight two directions for future work. Firstly, how does XRM relate to the invariance principle $Y\perp E\mid\Phi(X)$? What is the interplay between revealing relevant labels $Y$ and relevant environments $E$ as to afford invariance? To our knowledge, XRM is the first environment discovery algorithm tampering with labels $Y$, thus exploring invariance—and the violation thereof—from a new angle. Because relabeling happens with a probability proportional to confidence, we expect model calibration to play a role in understanding the theoretical underpinnings of XRM, as it happened with other invariance methods (Wald et al., 2021). Overall, the theoretical analysis of XRM will call for new tools, because label-flipping steers XRM away from the Bayes-optimal predictor.
Secondly, we would like to further understand the relationship between XRM and the multifarious phenomenon of memorization. Good memorization affords invariance (Where did I park my car?), and therefore depends on the collection of environments deemed relevant. Bad memorization happens due to “structured over-fitting”, commonly incarnated as a bad learning strategy; “use a simple feature for the majority, then memorize the minority”. XRM seems to attack a similar problem but, how does it specifically relate to these two flavours of memorization? Does XRM discover environments that promote features that benefit all examples?
@@ Acknowledgements
We thank Kartik Ahuja, Martín Arjovsky, Léon Bottou, Kamalika Chaudhuri, Cijo Jose, Reyhane Askari Hemmat and Polina Kirichenko for fruitful discussions. We thank the FAIR leadership for their help in publishing this manuscript.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Further Related Work
The literature on OOD and domain generalization spans a decade and comprises a vast amount of works. In the review below, we survey (i) some of the major milestones of OOD generalization research, (ii) advances in the sub-problem of sub-population shift, (iii) the multifarious connections between OOD generalization and causal inference, (iv) efforts to learn domains, sub-populations, or environments from pooled collections of training examples previous to our XRM proposal, and (v) their limitations in terms of annotation requirements and impossibility results.
  1. (i)
    The first works in OOD and domain generalization proposed algorithms that summarize each domain as a kernel mean embedding of the respective distribution of inputs (Blanchard et al., 2011; Muandet et al., 2013); these were later extended to the realm of deep neural networks (Zhang et al., 2021). One common avenue towards domain generalization is to learn a predictor where the feature representation has the same distribution across domains (Sun & Saenko, 2016; Ganin et al., 2016). Another major strategy is to enforce learning a richer feature space (Zhang et al., 2022a), which can be done by combining the weights of multiple models with different hyper-parameter configurations (Cha et al., 2021; Rame et al., 2022; Wortsman et al., 2022; Lin et al., 2023), or biasing training to make them disagree with each other (Nam et al., 2020; Pagliardini et al., 2023; Lee et al., 2023). Learning from combinations of examples, by means of mixup (Zhang et al., 2018), is also a promising route to diminish the impact of spurious correlations (Yao et al., 2022; Giannone et al., 2022). All in all, there are multiple frameworks that evaluate dozens of OOD generalization algorithms across a variety of benchmark datasets, such as DomainBed (Gulrajani & Lopez-Paz, 2020) and WILDS (Pang Wei Ko et al., 2021). We recommend the reader to consult recent surveys (Zhou et al., 2022a; Wang et al., 2021) for a taxonomy of the vast array algorithms on offer.
  2. (ii)
    Sub-population shift is a particular type of OOD generalization problem, where environments are direct annotations of a spurious attribute, and one can assume that the test domain will be equal a subdistribution—group—of the training data. The gold-standard for addressing sub-population shifts is group distributionally robust optimization (Sagawa et al., 2019, GroupDRO). Group subsampling and reweighting schemes, albeit simple, also provide state-of-the-art accuracy (Idrissi et al., 2022). To achieve good performance, it is known that it suffices to finetune the last layer of a deep neural network with a small training set with balanced groups (Izmailov et al., 2022). The framework of SubpopBench compares twenty algorithms for sub-population shift across a dozen benchmark datasets (Yang et al., 2023).
  3. (iii)
    The goal of OOD/domain generalization can be understood as finding predictors invariant across a family of relevant environments (Arjovsky et al., 2019). This establishes an intimate link between OOD generalization and causality under the interventionist account, where causation is defined as invariance across interventions (Woodward, 2005). A pioneering method attacks the problem OOD generalization as finding invariant causal predictors (Peters et al., 2016, ICP). The framework of invariant risk minimization (IRM) (Arjovsky et al., 2019) extends ICP to deep neural networks, advocating the invariance principle of “finding a feature representation such that the optimal classifier matches across environments”. Researchers have proposed multiple variants of the original IRM formulation, with notable examples being risk extrapolation (Krueger et al., 2021, vREX) and sparse risk minimization (Zhou et al., 2022b). The IRM framework has found multiple applications, with fair face recognition (Ma et al., 2023) being a recent example.
  4. (iv)
    The main factor limiting the application of OOD generalization and sub-population shift machinery is their requirement of domain, environment, or group annotations. Unfortunately, these are resource-intensive to obtain and are limited by human annotators’ biases, as the biases they identify may not align with those learned by models, and vice versa (Bell & Sagun, 2023).
    Consequently, a wide array of methods has been recently proposed to estimate these annotations from pooled collections of training data (LaBonte et al., 2023; Zhang et al., 2023; Tsirigotis et al., 2023). Among them, learning from failure (Nam et al., 2020, LfF) learns a biased network, and a final network that focuses on the examples misclassified by the biased network. Environment inference for invariant learning (Creager et al., 2020, EIIL) searches for an environmental partition that violates the IRM principle. Just-train-twice (Zheran Liu et al., 2021, JTT) trains one first network for a few iterations, and a final network to focus more on the examples from the first network. Correct and Contrast (Zhang et al., 2022b, CNC) leverages ERM failures and contrastive learning to learn a robust representation. Automatic feature reweighting (Qiu et al., 2023, AFR) learns a first network for a few iterations, and then fine-tunes the last layer to focus on mistakes. Learning to split (Bao & Barzilay, 2022, LS) and adversarial re-weighted learning (Lahoti et al., 2020, ARL) implement adversarial games to find a split of the training data inducing maximum out-of-sample error. Bias amplification (Li et al., 2023, BAM) incorporates per-example “slack variables” to absorb the fast learning of spurious correlations. No subclass left behind (Sohoni et al., 2020, GEORGE) clusters the hidden representation of a neural network to construct different environments. (Teney et al., 2021) manually identify variables to stratify pooled collections of training examples into environments. In the context of label noise robustness, the prior CrossSplit (Kim et al., 2023) uses a similar approach to that of XRM. CrossSplit uses a pair of networks trained on two disjoint parts of the training data and uses label correction to prevent memorization of noisy (mis-labelled) examples. While confidently mistaken cross-predictions are indicators of a model’s memorization of a noisy label in the context of the CrossSplit paper, they indicate a model’s reliance on spurious correlations in our work. In the context of graph OOD generalization, GALA (Chen et al., 2024) was proposed as a framework for graph OOD generalization that uses an assistant ERM model to identify invariant subgraphs through agreement and disagreement with proxy predictions.
  5. (v)
    One important note to the environment discovery methods described above is that they still require group annotations in a validation set, used for selecting a model with good worst-group-accuracy. In the complete absence of environment annotations, learning invariant predictors is an impossible task in its full generality (Chen et al., 2022; Lin et al., 2022a; Tan et al., 2023; Chen et al., 2023). Because we have proposed XRM as an alternative to surmount such daunting task, the next section provides intuitions to identify success and failure cases of our method.
Appendix B Experimental Details
B.1 Sub-population Shift Experiments
For the results in Section 4.1, we follow standard experimental protocol as in (Yang et al., 2023). For model selection, we adhere to the standard practice of using the worst-group-accuracy. We try 10 different hyper-parameter combinations detailed in Section B.7 with one random seed. We select the hyper-parameter combination and early-stopping iteration yielding maximal validation worst-group-accuracy (or, in the absence of groups, worst-class-accuracy). Next, we repeat the experiment 10 times with different seeds. Regardless of how training and validation groups are discovered, we always report test worst-group-accuracy over the human group annotations provided by each dataset.
We consider six standard datasets. These are the four image datasets Waterbirds (Wah et al., 2011), CelebA (Liu et al., 2015), MetaShift (Liang & Zou, 2022), and ImageNetBG (Xiao et al., 2020); and the two natural language datasets MultiNLI (Williams et al., 2017) and CivilComments (Borkan et al., 2019). For CelebA, predictors map pixel intensities into a binary “blonde/not-blonde” label. No individual face characteristics, landmarks, keypoints, facial mapping, metadata, or any other information was used to train our CelebA predictors. Image datasets use a pretrained ResNet-50 (He et al., 2016), frozen for XRM experiments. Text datasets use a pretrained BERT (Devlin et al., 2018). The linear layers on the top of the pretrained models are initialized at zero. All images are resized and center-cropped to $224\times 224$ pixels, and undergo no data augmentation. We use SGD with momentum $0.9$ to learn from image datasets unless otherwise mentioned, and we employ AdamW (Loshchilov & Hutter, 2017) with default $\beta_{1}=0.9$ and $\beta_{2}=0.999$ for text benchmarks. As for the OOD generalization algorithms, we consider ERM, group distributionally robust optimization (Sagawa et al., 2019, GroupDRO), group re-weighting (Japkowicz, 2000, RWG), and group sub-sampling (Idrissi et al., 2022, SUBG).
We also conduct experiments on ColorMNIST (Arjovsky et al., 2019), but keep a strict protocol. More specifically, we set both training and validation data to contain two environments, with $0.8$ and $0.9$ label-color correlation, while the test environment shows $0.1$ label-color correlation. This contrasts Arjovsky et al. (2019), who used the test environment for model selection. We train a three-layer fully-connected network with layer sizes $[2*14*14,300,300,2]$ and use ReLU as the activation function. The network is optimized using the Adam optimizer with a learning rate of $31e-$ 3, and default parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We train all algorithms for a number of iterations that allows convergence within a reasonable compute budget. Full results with error bars are reported in Section B.1.
Table 5: Worst-group-accuracies and standard deviations for all datasets, algorithms, and annotations over 10 trials.
@@ B.1 Sub-population Shift Experiments
Waterbirds 66.4 $\pm$ 0.9 66.4 $\pm$ 0.9 66.4 $\pm$ 0.9 67.3 $\pm$ 0.6 86.5 $\pm$ 0.5 88.1 $\pm$ 0.9 66.8 $\pm$ 0.5 87.3 $\pm$ 0.3 82.0 $\pm$ 1.0 62.1 $\pm$ 1.5 87.1 $\pm$ 1.1 77.7 $\pm$ 2.7
CelebA 54.3 $\pm$ 2.7 55.1 $\pm$ 1.8 58.6 $\pm$ 3.8 68.4 $\pm$ 2.1 88.3 $\pm$ 2.1 89.1 $\pm$ 1.3 64.7 $\pm$ 1.1 82.6 $\pm$ 1.8 81.1 $\pm$ 2.1 65.9 $\pm$ 2.7 83.9 $\pm$ 2.8 81.1 $\pm$ 2.3
MultiNLI 67.9 $\pm$ 2.0 72.0 $\pm$ 1.2 69.1 $\pm$ 1.9 68.6 $\pm$ 1.4 73.4 $\pm$ 4.8 72.1 $\pm$ 1.0 66.9 $\pm$ 1.8 71.3 $\pm$ 1.7 70.2 $\pm$ 1.3 69.7 $\pm$ 1.4 52.4 $\pm$ 25.8 72.0 $\pm$ 1.2
CivilComments 67.2 $\pm$ 1.3 74.0 $\pm$ 0.5 64.7 $\pm$ 9.8 66.6 $\pm$ 1.1 73.8 $\pm$ 0.6 72.2 $\pm$ 0.8 67.2 $\pm$ 1.3 73.4 $\pm$ 0.5 72.2 $\pm$ 0.4 65.4 $\pm$ 1.9 71.1 $\pm$ 1.4 44.6 $\pm$ 4.2
ColorMNIST 10.0 $\pm$ 0.0 10.1 $\pm$ 0.3 11.3 $\pm$ 0.9 9.9 $\pm$ 0.0 10.1 $\pm$ 0.2 69.7 $\pm$ 1.1 10.0 $\pm$ 0.0 10.3 $\pm$ 0.4 69.4 $\pm$ 1.8 10.0 $\pm$ 0.0 9.9 $\pm$ 0.3 65.2 $\pm$ 2.2
MetaShift 64.2 $\pm$ 1.0 69.8 $\pm$ 3.3 71.7 $\pm$ 2.1 72.5 $\pm$ 3.0 80.3 $\pm$ 0.9 77.5 $\pm$ 1.0 72.3 $\pm$ 2.5 76.3 $\pm$ 2.1 74.2 $\pm$ 2.0 69.7 $\pm$ 3.7 77.0 $\pm$ 3.4 77.6 $\pm$ 1.4
ImagenetBG 78.4 $\pm$ 1.2 78.4 $\pm$ 1.3 79.0 $\pm$ 0.8 78.3 $\pm$ 0.9 78.2 $\pm$ 0.9 78.2 $\pm$ 1.2 78.3 $\pm$ 1.5 79.0 $\pm$ 1.5 78.8 $\pm$ 1.4 79.5 $\pm$ 1.1 78.8 $\pm$ 1.2 77.2 $\pm$ 1.2
For the experiment on CIFAR-10 (Krizhevsky et al., 2009), we train a VGG-16 model (Simonyan & Zisserman, 2014) using SGD with a learning rate of $21e-$ 2 and a momentum of $0.9$.
B.2 DomainBed Experiments
For the results in Section 5.2 and Section B.2, we adhere to the original codebase from DomainBed (Gulrajani & Lopez-Paz, 2020). Each dataset is used to train a model on all but one environment, which is held out as the test domain. This process is repeated for each possible environment as the test domain, resulting in multiple training and testing splits for each dataset. We experiment with ERM and the CORAL algorithm as it is the best performing single-model (non-ensembling) method according to the DomainBed suite. For each triplet of (dataset, method, test environment), we sweep over 10 different hyper-parameter combinations detailed in Section B.7. We perform model selection based on the average accuracy over the validation environments, which is referred to as the ‘training domain validation set’ in the DomainBed paper. In Section B.2, we report the results for each environment when selected as the test environment. Additionally, we report the average and worst environment test accuracies. In settings without environment annotations (i.e., ERM), we combine all training environments and then split into one training and one validation set. In those cases with annotations, whether human-annotated or discovered by XRM, each training environment is divided into as many training and validation sets as the number of environments.
Table 6: Full results for the DomainBed suite.
{NiceTabular}
[…]
ERM (None) 96.70 64.85 74.20 76.15 77.97 64.85
CORAL (Human) 97.35 65.00 72.80 76.35 77.87 65.00
CORAL (XRM) 95.55 66.15 72.45 76.50 77.66 66.15
{NiceTabular}
[…]
ERM (None) 84.65 80.65 95.55 72.55 83.35 72.55
CORAL (Human) 84.90 80.75 96.60 77.70 84.98 77.70
CORAL (XRM) 81.90 77.30 96.90 79.15 83.81 77.30
{NiceTabular}
[…]
ERM (None) 59.50 52.25 74.15 76.00 65.47 52.25
CORAL (Human) 64.00 53.55 76.15 77.25 67.73 53.55
CORAL (XRM) 61.80 53.90 74.85 77.50 67.01 53.90
{NiceTabular}
[…]
ERM (None) 54.80 42.30 56.40 34.60 47.02 34.60
CORAL (Human) 58.20 39.25 59.45 37.15 48.51 37.15
CORAL (XRM) 59.20 45.10 56.10 38.00 49.60 38.00
{NiceTabular}
[…]
ERM (None) 47.40 14.75 37.45 9.30 42.10 39.15 31.69 9.30
CORAL (Human) 60.05 20.25 47.90 13.25 59.95 50.45 41.97 13.25
CORAL (XRM) 50.40 16.80 42.30 11.60 50.40 43.70 35.87 11.60
B.3 Dominoes Experiments
We conduct an additional experiment on the dataset of Dominoes (Geirhos et al., 2020). Dominoes is an image dataset in which images from two other datasets are concatenated. Particularly, the top half of an image shows MNIST digits from classes {0, 1}. The bottom half shows images from classes {coat, dress} of FashionMNIST in the MF (MNIST-MNISTFashion) version. And in the MC (MNIST-CIFAR) version, the bottom half shows images from classes {car, truck} of CIFAR-10. We followed a setup as described in (Kirichenko et al., 2022) (except that we used a smaller ConvNet) where the spurious MNIST part is strongly correlated (99% of the time) with the labels in the training data but drops to 50% (random) in the validation and test sets. Meanwhile, the core feature (either FashionMNIST or CIFAR) is always fully correlated with the labels. We compare three methods on these datasets: the Vanilla ERM, GroupDRO) using the ground-truth annotations111annotations indicate whether the MNIST half matches the label or not, and GroupDRO with XRM-inferred annotations. The results, shown in Table 7, suggest that the group annotations inferred by XRM are as effective as the ground-truth annotations.
Figure 4: Logit scatter plots for models trained on Dominoes-MF. The x-axis shows the logits with the original training examples, while the y-axis displays the logits for the same training examples with the MNIST part removed. These plots help reveal the extent of reliance on the spurious feature (MNIST digit) versus the core feature (FashionMNIST or Cifar). a) Vanilla ERM Model: Strong reliance on the spurious MNIST digit feature and partial dependence on the core feature. b) XRM Model: Strong reliance on the spurious MNIST digit. c) GroupDRO Model with XRM-Inferred annotations: point mostly on the diagonal suggesting invariance to the spurious feature, focusing only on the core feature. d) GroupDRO with ground-truth annotations for reference. We highlight that for XRM, it is desirable to only rely on the spurious feature since this will then enable the subsequent GroupDRO to learn the invariant core feature.
Table 7: Worst-group accuracy of three methods on two Dominoes datasets. Dominoes MF and MC are concatenation of MNIST digits with FashionMNIST and Cifar-10, respectively. The MNIST digit is spurious with 99% correlation in the training set and random correlation in the validation and test sets. The core feature (either FashionMNIST or Cifar) is however always fully correlated with the labels. GT denotes ground-truth group annotations.
ERM XRM+GroupDRO GT+GroupDRO
Dominoes MF 50.74 86.68 85.28
Dominoes MC 48.30 68.78 69.43
To further compare these methods, in Figure 4, we provide a scatter plot of the model logits for each method. Specifically, we feed two versions of all training examples to a fully trained model and store two sets of logits. One set is the result of the original training examples, and the other set is obtained with the MNIST part removed (replaced by average).
  • For an ERM model, the logits vary mostly along the x-axis and less along the y-axis, suggesting that the model is mostly relying on the spurious feature but also partially on the core feature.
  • An XRM model varies mostly along the x-axis and is almost invariant along the y-axis, indicating that the model strongly relies on the spurious feature alone and hence can effectively split the data into two environments according to the spurious correlation.
  • For a model trained with GroupDRO on XRM-annotations, points are almost all on the diagonal, suggesting that removing the spurious feature does not change the logits much, implying that the model is invariant to the spurious feature and only relies on the core feature.
  • A model trained with GroupDRO using the ground-truth annotations shows similar results, proving it mainly relies on the core feature.
B.4 Hyperparameter Sensitivity Analysis of XRM on Waterbirds
Here, we conduct a hyperparameter sensitivity experiment on the Waterbirds dataset and report the results in Figure 5. We fix the batch-size at 512 and vary the learning rate and weight decay coefficient to see how XRM performs with different hyperparameters. This grid of outcomes across different learning rate and weight decay settings shows that XRM works well across a wide range of hyperparameter combinations.
For each combination of learning rate and weight decay (each cell in the grid), we fully train an XRM model, get the group annotations, and then use GroupDRO with these inferred annotations. The hyperparameters for GroupDRO kept the same throughout all experiments. Hence, any variability in performance is only attributed to the XRM phase.
Figure 5: Hyperparameter sensitivity analysis of XRM on the Waterbirds. Here we evaluate the sensitivity of XRM’s performance to variations in learning rate and weight decay hyperparameters, while keeping the batch size fixed at 512. Left: worst-group-acc of a GroupDRO model trained with XRM annotations. Each cell within the grid represents a hyperparameter combination for XRM, with the color intensity indicating the test worst-group-accuracy. Right: the percentage of labels flipped by XRM for the corresponding hyperparameter combination, serving as the model selection criterion. Analysis: The best XRM model is chosen according to the highest flip percentage (red square), which resulted in an accuracy of 87.3% (dashed red square). However, the plot on the left shows that even neighboring cells with very different hyperparameters can still lead to near-optimal performance.
B.5 Can XRM Handle Settings with More Than Two True Underlying Environments?
XRM is always presented with pooled data with no explicit annotations. Regardless of how many human-annotation environments are in the pooled data, XRM always splits the data into 2 environments (2 groups per label). For example, all the datasets in table 6, have more than two human-annotated environments. That is while XRM split the examples into two environments.
Whether two environments suffice—when constructed appropriately—is in fact a fascinating question that we have pondered about for a while. Our experiments, especially those on the DomainBed, show that although XRM splits the data into a smaller number of groups than the human annotations, it is successful in forming sufficiently varied groups to match the performance with human annotations.
To illustrate, take one of the experiments with the PACS dataset where the “Photo” environment is left for testing and the other three environments of “Art”, “Cartoon”, and “Sketch” are pooled together and presented to XRM. Now XRM’s objective is to split the dataset into environments that lead to learning an invariant predictor, not to recover the original environment. XRM’s splitting is done according to the ‘hardness’ of examples and might or might not align with the human annotations. To see that, we visualize a per-class confusion matrix for the environment discovery of XRM. We also quantify the alignment between XRM environments and human-annotated ones using Normalized-Mutual-Information (NMI), a higher NMI means more alignment.
We observe that for some of the classes (e.g., the “House” class) there is a rather strong alignment. However, for some other classes (e.g., the “Dog” class), there is no alignment between human-annotated environments and those of XRM. That is absolutely fine as long as the inferred environments can be used for invariant learning and that appears to be the case.
As a reference point, a random splitter would lead to an NMI of zero on all classes and a subsequent invariant learning method cannot do better than an ERM with no environment annotations.
[…]
B.6 XRM Model Selection
To determine the best hyper-parameter combination for each experiment, we run XRM with 10 different hyper-parameter combinations detailed in Section B.7 and one random seed. We then compute the percentage of flipped labels that appear at the last iteration for each combination. The combination with the highest percentage is selected as the best. We disregard rare cases in which a hyper-parameter combination yields degenerate grouping where every example in one class in misclassified. Next, we repeat the experiment 10 times with different seeds to obtain 10 sets of inferred labels. Finally, we apply an OOD generalization algorithm on these inferred labels again using 10 different seeds.
B.7 Hyper-parameter Sampling Grids
algorithm hyper-parameter ResNet BERT
learning rate $10^{\text{Uniform}(-5,-3)}$ $10^{\text{Uniform}(-6,-4)}$
SUBG, RWG batch size $2^{\text{Uniform}(5,7)}$ $2^{\text{Uniform}(4,6)}$
dropout $\text{Random}([0,0.1,0.5])$
GroupDRO $\eta$ $10^{\text{Uniform}(-3,-1)}$ $10^{\text{Uniform}(-3,-1)}$
B.8 Statistics of the Datasets
Dataset Data type Number of envs. Number of classes Dataset size
Waterbirds Image 2 2 11788
CelebA Image 2 2 202599
CivilComments Text 8 2 242436
MultiNLI Text 2 3 412349
MetaShift Image 2 2 3499
ImageNetBG Image 2 9 192255
VLCS Image 4 5 10729
PACS Image 4 7 9991
OfficeHome Image 4 65 15588
Terra Incognita Image 4 10 24788
DomainNet Image 6 345 586575
B.9 Learning to Split on Waterbirds
We benchmarked the official learning to split code-base https://github.com/YujiaBao/ls on the WaterBirds dataset. We found on a Volta-32GB GPU running the learning to split group inference module took approximately 20 hours. We assessed the method’s sensitivity to two hyperparameters: the number of epochs used for early stopping (patience argument in the codebase) and the pre-supposed ratio of groups (based on the ratio argument in the code). For patience we swept over (2, 5, 10) with 5 being the default value. For ratio, we swept over (0.25, 0.5, 0.75) with 0.75 being the default value based on the paper. We found worst group performance using a fixed GroupDRO phase-2 training varied by as much as ±$\pm 7\%{}$ on Waterbirds.
Appendix C XRM in PyTorch